# soft open fences

### v2

@@ Abstract
We propose (S)peculative (C)ontrastive MCTS: a novel Monte Carlo Tree Search (MCTS) reasoning algorithm for Large Language Models (LLMs), significantly improves both reasoning accuracy and speed. Our motivation comes from: 1. Previous MCTS LLM reasoning works often overlooked its biggest drawback—slower speed compared to CoT; 2. Previous research mainly used MCTS as a tool for LLM reasoning on various tasks with limited quantitative analysis or ablation studies of its components from reasoning interpretability perspective. 3. The reward model is the most crucial component in MCTS, however previous work has rarely conducted in-depth study or improvement of MCTS’s reward models. Thus, we conducted extensive ablation studies and quantitative analysis on components of MCTS, revealing the impact of each component on the MCTS reasoning performance of LLMs. Building on this, (i) we designed a highly interpretable reward model based on the principle of contrastive decoding and (ii) achieved an average speed improvement of 51.9% per node using speculative decoding. Additionally, (iii) we improved UCT node selection strategy and backpropagation used in previous works, resulting in significant performance improvement. We outperformed o1-mini by an average of 17.4% on the Blocksworld multi-step reasoning dataset using Llama-3.1-70B with SC-MCTS. Our code is available at https://github.com/zitian-gao/SC-MCTS.
@@ 1 Introduction
With the remarkable development of Large Language Models (LLMs), models such as o1 (OpenAI, 2024a) have now gained strong ability to multi-step reasoning across complex tasks and solve problems that are more difficult than previous scientific, code, and mathematical problems. The reasoning task has long been considered challenging for LLMs. These tasks require converting a problem into a series of reasoning steps and then executing those steps to arrive at the correct answer. Recently, LLMs have shown great potential in addressing such problems. A key approach is using Chain of Thought (CoT) (Wei et al., 2024), where LLMs break down the solution into a series of reasoning steps before arriving at the final answer. Despite the impressive capabilities of CoT-based LLMs, they still face challenges when solving problems with an increasing number of reasoning steps due to the curse of autoregressive decoding (Sprague et al., 2024). Previous work has explored reasoning through the use of heuristic reasoning algorithms. For example, Yao et al. (2024) applied heuristic-based search, such as Depth-First Search (DFS) to derive better reasoning paths. Similarly, Hao et al. (2023) employed MCTS to iteratively enhance reasoning step by step toward the goal.
@@
To address the these challenges, we first went beyond previous LLM reasoning works who primarily treated MCTS as an tool rather than analyzing and improving its components, to study each component of MCTS related to the reward model. We redesigned the reward model for MCTS reasoning based on the principle of contrastive decoding: specifically, we first combined multiple highly interpretable reward functions in contrast with a small model, then identified their modes by clustering the prior distributions, and normalized the rewards separately for each mode. We found the Upper Confidence Bound applied on Trees (UCT) strategy that is often set a priori in most previous works failed to function in our experiments, and the performance is sensitive to choices of the exploration constant. We also refined the backpropagation of MCTS to prefer more steadily improving paths. These changes boost the performance of our MCTS system. Finally, we incorporated speculative decoding into MCTS to address the speed challenge.
The contributions of our new method, SC-MCTS, are summarized as follows:
@@
    We designed a novel, general action-level reward model for MCTS multi-step reasoning based on the principle of contrastive decoding, which requires no external tools, training, or datasets. Additionally, we found that previous works often failed to effectively harness multiple reward models, thus we proposed a statistical linear combination method. At the same time, we introduced token-level speculative decoding which incurs no additional cost and speeds up MCTS reasoning by an average of 52%, as contrastive decoding also requires a smaller language model.
@@ 3.3 Contrastive Decoding
We discuss vanilla Contrastive Decoding (CD) from Li et al. (2023), which improves text generation in LLMs by reducing errors like repetition and self-contradiction. CD uses the differences between an expert model and an amateur model, enhancing the expert’s strengths and suppressing the amateur’s weaknesses. The CD objective is defined as:
@@
where $p_{\text{EXP}}$ and $p_{\text{AMA}}$ are the expert and amateur probability distributions.
To avoid penalizing correct behavior of the amateur or promoting implausible tokens, CD applies an adaptive plausibility constraint using an $\alpha$-mask, which filters tokens by their logits against a threshold:
@@
Final logits are adjusted with a coefficient $(1+\beta)$, modifying the contrastive effect on output scores (Liu et al., 2021):
@@
3.4 Speculative Decoding
Based on vanilla Speculative Decoding (Leviathan et al., 2023), the process can be summarized as follows: Let $M_{p}$ be the target model we aim to accelerate, with $p(x_{t}|x_{<t})$ representing its distribution for a given prefix $x_{<t}$. Let $M_{q}$ be a smaller, more efficient approximation model, with $q(x_{t}|x_{<t})$ representing its distribution. The key idea is to generate $\gamma$ new tokens using $M_{q}$ and evaluate them against $M_{p}$, accepting those that align with $M_{p}$’s distribution. Each evaluation of $M_{p}$ on the new tokens which is parallel can produce at least one new token according to $p(x)$ in the end, with potentially many more tokens if they get accepted.
Specifically, speculative decoding first samples $x\sim q(x)$ autoregressively $\gamma$ times and keep them as long as $q(x)\leq p(x)$. If $q(x)>p(x)$ at some point, that sample is rejected with probability $1-\frac{p(x)}{q(x)}$, and a new sample is drawn from an adjusted distribution:
$p^{\prime}(x)=\text{norm}(\max(0,p(x)-q(x)))$
This ensures that the final sample distribution always follows $p(x)$. Since contrastive decoding and speculative decoding both require smaller language model, after employing our proposed action level contrastive decoding, we can achieve the acceleration effect of speculative decoding without additional cost (Yuan et al., 2024a).
@@ 4 Method
The SC-MCTS reasoning methodology is outlined as follows: In Section 4.1 details the multi-reward design; Section 4.2 discusses the UCT strategy for node selection; and Section 4.3 proposes the refinement of backpropagation.
@@ 4.1 Multi-Reward Design
5:     $\bm{\mu}_{r},\bm{\sigma}_{r}\leftarrow\text{Cluster-stats}(r(\tilde{A}),K)$triangleright Prior statistics (Equation 1)
@@
Inspired by contrastive decoding, we propose a novel reward model: average Jensen-Shannon divergence between the expert model’s logits and the amateur model’s logits. Unlike vanilla contrastive decoding (Li et al., 2023), which operates at the token level, our novel reward is computed at the action level, treating a sequence of action tokens as a whole:
@@
where $p_{\text{e}}$ and $p_{\text{a}}$ represent the softmax probabilities of the expert model and the amateur model, respectively. This approach ensures that the reward captures model behavior at the action level, as the entire sequence of tokens is taken into account at once. This contrasts with token-level methods like vanilla contrastive decoding, where each token is treated serially.
@@
After generating the answer, we prompt the model to evaluate the quality of its response by asking "Is this answer correct/good?" This serves to capture the model’s confidence in its own output, leading to more informed decision-making. The self evaluation prompt’s logits are then used to calculate a reward function. Similar to the loglikelihood reward model, we calculate the self evaluation reward $R_{\text{SE}}$ by summing the log-probabilities over the self-evaluation tokens or obtaining the log-probability of a single ’good’ token in between.
Harnessing Multiple Reward Models and Modes
We collected prior distributions for the reward factors and found some of them span multiple regions. Therefore, we compute the fine-grained prior statistics as mean and standard deviation of modes of the prior distribution ${\mathcal{R}}\in\{{\mathcal{R}}_{\text{JSD}},{\mathcal{R}}_{\text{LL}},{\mathcal{R}}_{\text{SE}}\}$:
@@
This reward design, which we call Multi-RM method, has some caveats: first, to prevent distribution shift during reasoning, we update the mean and standard deviation of the reward functions online for each mode (see Appendix D for pseudocode); second, we focus only on cases with clearly distinct reward modes, leaving general cases for future work. For the correlation heatmap, see Appendix E.
@@ 5.2 Main Results
To evaluate the SC-MCTS algorithm in LLM multi-step reasoning, we implemented CoT, RAP-MCTS, and SC-MCTS using Llama-3-70B and Llama-3.1-70B. For comparison, we used Llama-3.1-405B and GPT-4o for CoT, and applied 0 and 4 shot single turn for o1-mini, as OpenAI (2024b) suggests avoiding CoT prompting. The experiment was conducted on Blocksworld dataset across all steps and difficulties. For LLM settings, GPU and OpenAI API usage data, see Appendix C and G.
@@ 5.4 Parameters
As discussed in Section 4.2, the constant $C$ is a crucial part of UCT strategy, which completely determines whether the exploration term takes effect. Therefore, we conducted quantitative experiments on the constant $C$, to eliminate interference from other factors, we only use MCTS base with the common reward model $R_{\text{LL}}$ for both RAP-MCTS and SC-MCTS. From Figure 5 we can observe that the constant $C$ of RAP-MCTS is too small to function effectively, while the constant $C$ of SC-MCTS is the value most suited to the values of reward model derived from extensive experimental data.
@@ 5.6 Interpretability Study
Figure 6: Distribution of reward functions of SC-MCTS on the Blocksworld multi-step reasoning dataset using Llama-3.1-70B as expert model and Llama-3.2-1B as amateur model.
We observe two distinct types of reward value distributions: normal distribution and half-normal distribution (truncated normal distribution).
Normal distribution is represented by $R_{\text{LL}}$ which is computed by accumulating the log probabilities over the token indices of the answer portion from the logits. This accumulation captures the probability trends of the entire answer generation, with a symmetric and concentrated distribution, resulting in a normal distribution. This distribution reflects the balance in the reasoning process and provides a stable and interpretable guide to MCTS multi-step reasoning.
Half-normal distribution includes $R_{\text{JSD}}$ and $R_{\text{SE}}$. $R_{\text{JSD}}$ quantifies the difference in confidence between the expert model and the amateur model over the logits of the answer part. The motivation for designing this reward function stems from the principle of contrastive decoding where constraining the expert model with the amateur model significantly reduces the low-quality outputs of the expert (Li et al., 2023). $R_{\text{JSD}}$ introduces symmetry by averaging distributions defined as $M=\frac{1}{2}(\text{Logits}_{\text{Expert}}+\text{Logits}_{\text{Amateur}})$, which differs from KL divergence. This symmetry combined with the non-negative nature of JS divergence also leads to a half-normal distribution. $R_{\text{JSD}}$ guide MCTS reasoning by measuring the discrepancy between the models’ logits. Our ablation experiments in Section 5.5 demonstrate that $R_{\text{JSD}}$ is highly effective as a reward function, significantly improving the performance of multi-step reasoning.
The self-evaluation reward function computes the log probability of a single self evaluation token such as "good" to measure the model’s confidence in generated answer (Ren et al., 2023). Since it only focuses on the log probability of single specific token and relies on the model’s confidence in particular outputs, its distribution also takes the form of a half-normal distribution. This distribution reflects the uncertainty in the reasoning process especially at decision points with lower confidence.
By combining the observations from ablation experiments in Table 2 and Figure 6, we also find that reward functions whose distributions more closely resemble a normal or half-normal distribution tend to perform better. For example, $R_{\text{JSD}}$, whose shape is more aligned with a half-normal distribution compared to $R_{\text{SE}}$, also demonstrates better performance than $R_{\text{SE}}$.
From the ablation study results in Table 5.5, it can be seen that the reward model is the most critical factor for MCTS reasoning performance. Having well-interpretable reward models implies better interpretability of MCTS reasoning. By studying these highly interpretable rewards, we can gain a clearer and more explainable understanding of MCTS reasoning.
@@ 6 Conclusion
In this paper, we present SC-MCTS, a novel and effective algorithm to enhancing the reasoning capabilities of LLMs. With extensive improvements in reward modeling, node selection strategy and backpropagation, SC-MCTS boosts both accuracy and speed, outperforming OpenAI’s o1-mini model by 17.4% on average using Llama-3.1-70B on the Blocksworld dataset. Experiments demonstrate its strong performance, making it a promising approach for multi-step reasoning tasks. For future work please refer to Appendix I. The synthesis of interpretability, efficiency and generalizability positions SC-MCTS as a valuable contribution to advancing LLMs multi-step reasoning.

### v3

@@ Abstract
We propose (S)peculative (C)ontrastive MCTS: a novel Monte Carlo Tree Search (MCTS) reasoning algorithm for Large Language Models (LLMs) which significantly improves both reasoning accuracy and speed. Our motivation comes from: 1. Previous MCTS LLM reasoning works often overlooked its biggest drawback—slower speed compared to CoT; 2. Previous research mainly used MCTS as a tool for LLM reasoning on various tasks with limited quantitative analysis or ablation studies of its components from reasoning interpretability perspective. 3. The reward model is the most crucial component in MCTS, however previous work has rarely conducted in-depth study or improvement of MCTS’s reward models. Thus, we conducted extensive ablation studies and quantitative analysis on components of MCTS, revealing the impact of each component on the MCTS reasoning performance of LLMs. Building on this, (i) we designed a highly interpretable reward model based on the principle of contrastive decoding and (ii) achieved an average speed improvement of 51.9% per node using speculative decoding. Additionally, (iii) we improved UCT node selection strategy and backpropagation used in previous works, resulting in significant performance improvement. We outperformed o1-mini by an average of 17.4% on the Blocksworld multi-step reasoning dataset using Llama-3.1-70B with SC-MCTS. Our code is available at https://github.com/zitian-gao/SC-MCTS.
@@ 1 Introduction
With the remarkable development of Large Language Models (LLMs), models such as o1 (OpenAI, 2024a) have now gained a strong ability for multi-step reasoning across complex tasks and can solve problems that are more difficult than previous scientific, code, and mathematical problems. The reasoning task has long been considered challenging for LLMs. These tasks require converting a problem into a series of reasoning steps and then executing those steps to arrive at the correct answer. Recently, LLMs have shown great potential in addressing such problems. A key approach is using Chain of Thought (CoT) (Wei et al., 2024), where LLMs break down the solution into a series of reasoning steps before arriving at the final answer. Despite the impressive capabilities of CoT-based LLMs, they still face challenges when solving problems with an increasing number of reasoning steps due to the curse of autoregressive decoding (Sprague et al., 2024). Previous work has explored reasoning through the use of heuristic reasoning algorithms. For example, Yao et al. (2024) applied heuristic-based search, such as Depth-First Search (DFS) to derive better reasoning paths. Similarly, Hao et al. (2023) employed MCTS to iteratively enhance reasoning step by step toward the goal.
@@
To address the these challenges, we went beyond prior works that treated MCTS as a tool and focused on analyzing and improving its components especially reward model. Using contrastive decoding, we redesigned reward model by integrating interpretable reward signals, clustering their prior distributions, and normalizing the rewards using our proposed prior statistical method. To prevent distribution shift, we also incorporated an online incremental update algorithm. We found that the commonly used Upper Confidence Bound on Trees (UCT) strategy often underperformed due to sensitivity to the exploration constant, so we refined it and improved backpropagation to favor steadily improving paths. To address speed issues, we integrated speculative decoding as a "free lunch." All experiments were conducted using the Blocksworld dataset detailed in Section 5.1.
Our goal is to: (i) design novel and high-performance reward models and maximize the performance of reward model combinations, (ii) analyze and optimize the performance of various MCTS components, (iii) enhance the interpretability of MCTS reasoning, (iv) and accelerate MCTS reasoning. Our contributions are summarized as follows:
@@
    We designed a novel, general action-level reward model based on the principle of contrastive decoding, which requires no external tools, training, or datasets. Additionally, we found that previous works often failed to effectively harness multiple reward models, thus we proposed a statistical linear combination method. At the same time, we introduced speculative decoding to speed up MCTS reasoning by an average of 52% as a "free lunch."
@@ 3.3 Contrastive Decoding
We discuss vanilla Contrastive Decoding (CD) from Li et al. (2023), which improves text generation in LLMs by reducing errors like repetition and self-contradiction. CD uses the differences between an expert model and an amateur model, enhancing the expert’s strengths and suppressing the amateur’s weaknesses. Consider a prompt of length $n$, the CD objective is defined as:
@@
where $x_{\text{pre}}$ is the sequence of tokens $x_{1},\dots,x_{n}$, the model generates continuations of length $m$, $x_{\text{cont}}$ is the sequence of tokens $x_{n+1},\dots,x_{n+m}$, and $p_{\text{EXP}}$ and $p_{\text{AMA}}$ are the expert and amateur probability distributions. To avoid penalizing correct behavior of the amateur or promoting implausible tokens, CD applies an adaptive plausibility constraint using an $\alpha$-mask, which filters tokens by their logits against a threshold, the filtered vocabulary $V_{\text{valid}}$ is defined as:
@@
where $s^{(i)}_{\text{EXP}}$ and $s^{(i)}_{\text{AMA}}$ are unnormalized logits assigned to token i by the expert and amateur models. Final logits are adjusted with a coefficient $(1+\beta)$, modifying the contrastive effect on output scores (Liu et al., 2021):
@@
3.4 Speculative Decoding as "free lunch"
Based on Speculative Decoding (Leviathan et al., 2023), the process can be summarized as follows: Let $M_{p}$ be the target model with the conditional distribution $p(x_{t}|x_{<t})$, and $M_{q}$ be a smaller approximation model with $q(x_{t}|x_{<t})$. The key idea is to generate $\gamma$ tokens using $M_{q}$ and filter them against $M_{p}$’s distribution, accepting tokens consistent with $M_{p}$. Speculative decoding samples $\gamma$ tokens autoregressively from $M_{q}$, keeping those where $q(x)\leq p(x)$. If $q(x)>p(x)$, the sample is rejected with probability $1-\frac{p(x)}{q(x)}$, and a new sample is drawn from the adjusted distribution:
$p^{\prime}(x)=\text{norm}(\max(0,p(x)-q(x))).$
Since both contrastive and speculative decoding rely on the same smaller models, we can achieve the acceleration effect of speculative decoding as a "free lunch" (Yuan et al., 2024a).
@@ 4.1 Multi-Reward Design
Our primary goal is to design novel and and high-performance reward models for MCTS reasoning and to maximize the performance of reward model combinations, as our ablation experiments in Section 5.5 demonstrate that MCTS performance is almost entirely determined by the reward model.
@@
5:     $\bm{\mu}_{r},\bm{\sigma}_{r},\bm{b}_{r}\leftarrow\text{Cluster-stats}(r(\tilde{A}),K)$triangleright Prior statistics (Equation 1)
@@
The Jensen-Shannon divergence (JSD) is a symmetric and bounded measure of similarity between two probability distributions $P$ and $Q$. It is defined as:
$\mathrm{JSD}(P\,\|\,Q)=\frac{1}{2}\mathrm{KL}(P\,\|\,M)+\frac{1}{2}\mathrm{KL}(Q\,\|\,M),\quad M=\frac{1}{2}(P+Q),$
where $\mathrm{KL}(P\,\|\,Q)$ is the Kullback-Leibler Divergence (KLD), and $M$ represents the midpoint distribution. The JSD is bounded between 0 and 1 for discrete distributions, making it better than KLD for online normalization of reward modeling.
Inspired by contrastive decoding, we propose our novel reward model: JSD between the expert model’s logits and the amateur model’s logits. Unlike vanilla token-level contrastive decoding (Li et al., 2023), our reward is computed at action-level, treating a sequence of action tokens as a whole:
@@
where $n$ is the length of tokens, $T_{\text{prefix}}$ is the index of the last prefix token, $p_{\text{e}}$ and $p_{\text{a}}$ represent the softmax probabilities of the expert and amateur models, respectively. This approach ensures that the reward captures model behavior at the action level as the entire sequence of tokens is taken into account at once. This contrasts with vanilla token-level methods where each token is treated serially.
@@
After generating the answer, we prompt the model to self-evaluate its response by asking "Is this answer correct/good?" This serves to capture the model’s confidence in its own output leading to more informed decision-making. The self evaluation prompt’s logits are then used to calculate a reward function. Similar to the loglikelihood reward model, we calculate the self evaluation reward $R_{\text{SE}}$ by summing the log-probabilities over the self-evaluation tokens.
Harnessing Multiple Reward Models
We collected prior distributions for the reward models and found some of them span multiple regions. Therefore, we compute the fine-grained prior statistics as mean and standard deviation of modes of the prior distribution ${\mathcal{R}}\in\{{\mathcal{R}}_{\text{JSD}},{\mathcal{R}}_{\text{LL}},{\mathcal{R}}_{\text{SE}}\}$:
@@
This reward design, which we call Multi-RM method, has some caveats: first, to prevent distribution shift during reasoning, we update the mean and standard deviation of the reward functions online for each mode (see Appendix D for pseudocode); second, we focus only on cases with clearly distinct reward modes, leaving general cases for future work. For the correlation heatmap, see Appendix C.
@@ 5.2 Main Results
To evaluate the SC-MCTS algorithm in LLM multi-step reasoning, we implemented CoT, RAP-MCTS, and SC-MCTS using Llama-3-70B and Llama-3.1-70B. For comparison, we used Llama-3.1-405B and GPT-4o for CoT, and applied 0 and 4 shot single turn for o1-mini, as OpenAI (2024b) suggests avoiding CoT prompting. The experiment was conducted on Blocksworld dataset across all steps and difficulties. For LLM settings, GPU and OpenAI API usage data, see Appendix E and H.
@@ 5.4 Parameters
As discussed in Section 4.2, the constant $C$ is a crucial part of UCT strategy, which completely determines whether the exploration term takes effect. Therefore, we conducted quantitative experiments on the constant $C$, to eliminate interference from other factors, we only use MCTS base with the common reward model $R_{\text{LL}}$ for both RAP-MCTS and SC-MCTS. From Figure 5 we can observe that the constant $C$ of RAP-MCTS is too small to function effectively, while the constant $C$ of SC-MCTS is the value most suited to the values of reward model derived from extensive experimental data. After introducing new datasets, this hyperparameter may need to be re-tuned.
@@ 5.6 Interpretability Study
In the Blocksworld multi-step reasoning dataset, we utilize a built-in ground truth verifier to measure the percentage of progress toward achieving the goal at a given step, denoted as $P$. The value of $P$ ranges between $[0,1]$. For any arbitrary non-root node $N_{i}$, the progress is defined as:
$P(N_{i})=\text{Verifier}(N_{i}).$
For instance, in a 10-step Blocksworld reasoning task, the initial node $A$ has $P(A)=0$. After executing one correct action and transitioning to the next node $B$, the progress becomes $P(B)=0.1$.
Given a non-root node $N_{i}$, transitioning to its parent node $\text{Parent}(N_{i})$ through a specific action $a$, the contribution of $a$ toward the final goal state is defined as:
$\Delta_{a}=P(\text{Parent}(N_{i}))-P(N_{i}).$
Next, by analyzing the relationship between $\Delta_{a}$ and the reward value $R_{a}$ assigned by the reward model for action $a$, we aim to reveal how our designed reward model provides highly interpretable reward signals for the selection of each node in MCTS. We also compare the performance of our reward model against a baseline reward model. Specifically, the alignment between $\Delta_{a}$ and $R_{a}$ demonstrates the interpretability of the reward model in guiding the reasoning process toward the goal state. Since Section 5.5 has already demonstrated that the reasoning performance of MCTS reasoning is almost entirely determined by the reward model, using interpretable reward models greatly enhances the interpretability of our algorithm SC-MCTS.
Figure 6: Reward distribution and interpretability analysis. The left histogram shows the baseline reward model (RAP-MCTS), while the right represents SC-MCTS. Bin colors indicate the proportion of positive $\Delta_{a}$ (lighter colors means higher proportions). Spearman and Pearson correlations along with p-values are shown in the top right of each histogram.
From Figure 6, shows that SC-MCTS* reward values correlate significantly with $\Delta_{a}$, as indicated by the high Spearman and Pearson coefficients. Additionally, the mapping between the reward value bins and the proportion of positive $\Delta_{a}$ (indicated by the color gradient from light to dark) is highly consistent and intuitive. This strong alignment suggests that our reward model effectively captures the progress toward the goal state, providing interpretable signals for action selection during reasoning.
These results highlight the exceptional interpretability of our designed reward model, which ensures that SC-MCTS* not only achieves superior reasoning performance but is also highly interpretable. This interpretability is crucial for understanding and improving the decision-making process in multi-step reasoning tasks, further validating transparency of our proposed algorithm.
@@ 6 Conclusion
In this paper, we present SC-MCTS, a novel and effective algorithm to enhancing the reasoning capabilities of LLMs. With extensive improvements in reward modeling, node selection strategy and backpropagation, SC-MCTS boosts both accuracy and speed, outperforming OpenAI’s o1-mini model by 17.4% on average using Llama-3.1-70B on the Blocksworld dataset. Experiments demonstrate its strong performance, making it a promising approach for multi-step reasoning tasks. For future work please refer to Appendix J. The synthesis of interpretability, efficiency and generalizability positions SC-MCTS as a valuable contribution to advancing LLMs multi-step reasoning.